# SynopGround: A Large-Scale Dataset for Multi-Paragraph Video Grounding from TV Dramas and Synopses

### v1

@@ 1. Introduction
However, considerable drawbacks still exist in previous video grounding datasets (Gao et al., 2017; Anne Hendricks et al., 2017; Krishna et al., 2017; Regneri et al., 2013; Grauman et al., 2022; Soldan et al., 2022). First of all, as presented in Figure 1, most of commonly-used datasets are constructed upon short videos and brief sentence queries. This setup limits the model in developing stronger abilities that can model and bridge the long-form videos (Yang et al., 2020; Gao et al., 2017) and long-text queries (Hou et al., 2023; Pan et al., 2023). Besides, shorter queries that describe detailed events (as shown in Table 1), are more prone to causing the risk of semantic ambiguity in referring expressions (Mao et al., 2016; Otani et al., 2020), i.e., the occurrence of one-to-many correspondence between queries and moments, which will adversely affect the model learning. In particular, this ambiguity issue is more prominent for the recently proposed MAD (Soldan et al., 2022) dataset which features long input videos but short general descriptions. For example, it is highly likely to find cases where multiple moments are semantically corresponding to the same short description like “She steps closer.” (shown in Table 1), especially when searching content in a long video. Furthermore, as listed in Table 1, existing benchmarks are tailored for language queries referring to low-level visible activities, while all of them overlook the importance of more complex events and abstract concepts. Such drawbacks actually limit the applications of video grounding in scenarios where complex descriptions with abstract concepts should be associated with long-term videos. For example, accelerating the movie post-production by automatically integrating raw footage into a coherent story based on the plot scripts is a practical need, but it cannot be satisfied by the current video grounding techniques developed from existing datasets.
@@
  • We are the first to research video-language grounding involving long videos and long queries. Comparison results show the unique advantages of our dataset and the efficacy of our baseline in multi-paragraph video grounding.
@@ 2.2. Narrative Video Understanding
Understanding visual content presented in narrative videos is an important area with many works (Simoes et al., 2016; Haurilet et al., 2016; Nagrani and Zisserman, 2018; Huang et al., 2018; Gu et al., 2018; Del Fabro and Böszörmenyi, 2013; Huang et al., 2020; Rao et al., 2020; Tapaswi et al., 2016; Jasani et al., 2019; Soldan et al., 2022; Xiong et al., 2019; Bain et al., 2020; Sun et al., 2022) proposed accordingly. Most of these prior works neglect to model and understand the content of narrative videos based on their high-level storylines while focusing on specific downstream applications, such as movie genre classification (Simoes et al., 2016), character identification (Haurilet et al., 2016; Nagrani and Zisserman, 2018; Huang et al., 2018), action localization (Gu et al., 2018), scene segmentation (Del Fabro and Böszörmenyi, 2013; Huang et al., 2020), and shot classification (Rao et al., 2020). In addition to that, some works have started pursuing story-level understanding in many different ways. For instance, Tapaswi et al.  (Tapaswi et al., 2016) proposed MovieQA dataset to comprehend movie stories by question-answering. MSA (Xiong et al., 2019), CMD (Bain et al., 2020), and SyMoN (Sun et al., 2022) datasets utilized synopses as language queries and formulated movie understanding as text-to-video retrieval. This line of research is closely related to ours. However, we tackle a more challenging video grounding task with long contexts and complex queries, in which the model should understand the long-range cross-modal dependencies so as to reason about the video grounding results at story level.
@@ 3.1. Problem Formulation
Considering that previous video paragraph grounding (Bao et al., 2021) is limited to a multi-query version of short single sentence grounding, we introduce a more challenging multi-query video grounding setting called Multi-Paragraph Video Grounding (MPVG). Specifically, given an untrimmed video $\mathcal{V}$ and $N$ consecutive paragraph queries $\mathcal{Q}=\{Q_{1},Q_{2},...,Q_{N}\}$ as input, the output should be $N$ temporal intervals $\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{N}\}$ corresponding to each of the paragraph queries, where $\mathcal{T}_{i}=\left(t_{s}^{(i)},t_{e}^{(i)}\right)$ indicates the starting timestamp $t_{s}^{(i)}$ and ending timestamp $t_{e}^{(i)}$ for the $i$-th paragraph query in the target video. In our dataset, the video $\mathcal{V}$ is an episode from a TV drama, and $\mathcal{Q}$ is the corresponding human-written synopsis that contains $N$ paragraphs, with $Q_{i}$ indicating the $i$-th paragraph in the synopsis $\mathcal{Q}$. Illustration of the MPVG problem is in Figure 2.
@@ 3.2. Data Collection and Annotation
We collect all the TV drama episodes from a leading online platform Tencent Video with official acknowledgement and permission. The plot synopsis for each episode of the TV dramas is scraped from a specialized TV review website [^1] URL: https://www.tvmao.com. All texts are translated into English using Tencent Cloud Translation with full permission and compliance. that contains lots of synopses of the most popular TV drama episodes written by professionals. Synopses that are too long or too short are discarded to ensure an adequate number of paragraphs in each synopsis. A total of 520 licensed and high-viewership TV dramas with textual synopses are finally selected to constitute our dataset, and we randomly sample several episodes from each selected TV drama to further annotate. Specifically, annotators are asked to read and understand the synopsis in advance. They then thoroughly watch the corresponding TV drama episode to determine the starting and ending time of the video content depicted by each synopsis paragraph.
Our data annotation pipeline is organized into multiple rounds to ensure the annotation quality. Specifically, all collected videos are first divided into numerous disjoint subsets of videos. In each annotation round, synopsis paragraphs for videos in one subset will be annotated with timestamps and each annotator is told to provide a score to indicate level of confidence in the annotated results. Afterwards, we first discard samples with low confidence as an initial cleanup, and then some of the remaining samples are selected to be manually checked in terms of quality. If the annotation quality is thought of as satisfactory, the annotation process will move on to another unlabeled subset of video data. Otherwise, the current batch of data would be re-annotated. The above procedures are repeated by tens of annotators until we finish the annotation of all candidate samples. For post-annotation assessment, we randomly select a proportion of the annotated data to be re-annotated by other annotators. Concretely, we calculate the temporal IoU (Intersection over Union) between the two results from different annotators, which reaches a value of about 85%. This assessment result is much better than that of other datasets, such as the ActivityNet (Krishna et al., 2017), where different annotators only achieve an agreement degree around 70%. The higher degree of agreement across different annotators in our dataset verifies the effectiveness of our designed pipeline for data annotation and quality control.
@@ 3.3. Data Statistics
Detailed comparison with other datasets. In Table 2, we compare our dataset with other existing datasets in detail. As suggested, our videos are much longer in duration than those of Charades-STA (Gao et al., 2017), ActivityNet-Captions (Krishna et al., 2017), DiDeMo (Anne Hendricks et al., 2017), TACoS (Regneri et al., 2013) and Ego4d-NLQ (Grauman et al., 2022). Although the average video duration in our dataset is shorter than that of MAD, our total duration of videos is more than twice that of MAD, showing that our dataset is at a larger scale. Furthermore, the duration of target segments in our dataset is significantly longer while the normalized target span is still short, making our target moments challenging to be localized. Note that some datasets like MAD have shorter normalized target span than ours, but their short and general queries bring the harmful and undesirable semantic ambiguity issue as mentioned before. In addition, our dataset is the first benchmark to introduce paragraph queries, and the average number of words in each query is significantly larger than that of other datasets, which greatly reduces the semantic ambiguity of the queries. Moreover, our synopsis queries involve both abstract expressions and concrete descriptions, enabling the model to learn semantic concepts at more diverse abstraction levels.
Data splits. As shown by the statistics in Table 3, we carefully divide the entire data into three disjoint splits for training, validation, and testing. Our training, validation, and testing sets consist of 3,187, 400, and 400 videos, respectively. It is worth noting that each video is an episode from a TV drama. Additionally, we manually guarantee that half of the videos for validation/testing are sourced from TV dramas that are not selected for the training set.
@@ 4.4. Model Training
(6) $\mathcal{L}_{loc}$ $=\frac{1}{B}\sum_{i=1}^{B}\left[\frac{1}{\lambda_{1}}\mathcal{L}_{l1}(\mathcal{\hat{T}}_{i},\mathcal{T}_{i})+\mathcal{L}_{GIoU}(\mathcal{\hat{T}}_{i},\mathcal{T}_{i})\right],$
(7) $\mathcal{L}_{att}$ $=-\frac{1}{B}\sum_{i=1}^{B}\text{log}\left(\sum_{j=1}^{T}m_{ij}\cdot a_{ij}\right)$
where $B$ is the batch size. $\mathcal{L}_{l1}$ and $\mathcal{L}_{GIoU}$ are L1 and GIoU (Rezatofighi et al., 2019) losses, respectively. $\mathcal{\hat{T}}_{i}$ is the predicted time span for the $i$-th query and $\mathcal{T}_{i}$ is the ground-truth. $\mathcal{L}_{att}$ is an attention loss on the global query features. The term $a_{ij}$ indicates the attention weights between the $i$-th global query feature and the $j$-th video feature, while $m_{ij}$ is an indicator that takes 1 if the $j$-th video feature is inside the ground-truth interval of the $i$-th query, and 0 otherwise. This loss explicitly encourages the model to learn higher attention weights between text queries and visual elements that are correlated. In total, the training loss is defined as the weighted sum of the above two losses as $\mathcal{L}=\lambda_{1}\mathcal{L}_{loc}+\lambda_{2}\mathcal{L}_{att}$, where $\lambda_{1}$ and $\lambda_{2}$ are the hyper-parameters to balance these two different kinds of losses.
@@ 5.1. Experimental Setup
Implementation details. Our proposed method is implemented by PyTorch. The pre-extracted SlowFast, CLIP, and OCR features are aligned at sequence dimension and concatenated at channel dimension as the video feature input. We use a pre-trained RoBERTa (Liu et al., 2019) model to extract OCR features at each timestamp. The loss weights are set as $\lambda_{1}=1,\lambda_{2}=0.2$. For data augmentation, we choose to randomly shuffle the order of paragraphs in the same synopsis by a probability of $p$ during training. $p$ is determined by $\alpha=1-\frac{T_{i}}{T_{max}}$, where $T_{i}$ is the index of the current training epoch and $T_{max}$ is set to 20. We adopt a local window length $M$ of 25 for our video encoder. The number of layers for video encoder and query decoder are set as 2 and 3, respectively. Our model is trained on 4 NVIDIA Tesla V100 GPUs by Adam optimizer using a learning rate of 0.0001 and batch size of 16 for a total of 50 epochs within one day.
@@ Acknowledgements.
This work is supported by Guangdong Provincial Key Laboratory of Information Security Technology (No.2023B1212060026).

### v3

@@ 1. Introduction
However, considerable drawbacks still exist in previous video grounding datasets (Gao et al., 2017; Anne Hendricks et al., 2017; Krishna et al., 2017; Regneri et al., 2013; Grauman et al., 2022; Soldan et al., 2022). First of all, as presented in Figure 1, most of commonly-used datasets are constructed upon short videos and brief sentence queries. This setup limits the model in developing stronger abilities that can model and bridge the long-form videos (Hou et al., 2023; Pan et al., 2023) and long-text queries (Yang et al., 2020; Gao et al., 2017). Besides, shorter queries that describe detailed events (as shown in Table 1), are more prone to causing the risk of semantic ambiguity in referring expressions (Mao et al., 2016; Otani et al., 2020), i.e., the occurrence of one-to-many correspondence between queries and moments, which will adversely affect the model learning. In particular, this ambiguity issue is more prominent for the recently proposed MAD (Soldan et al., 2022) dataset which features long input videos but short general descriptions. For example, it is highly likely to find cases where multiple moments are semantically corresponding to the same short description like “She steps closer.” (shown in Table 1), especially when searching content in a long video. Furthermore, as listed in Table 1, existing benchmarks are tailored for language queries referring to low-level visible activities, while all of them overlook the importance of more complex events and abstract concepts. Such drawbacks actually limit the applications of video grounding in scenarios where complex descriptions with abstract concepts should be associated with long-term videos. For example, accelerating the movie post-production by automatically integrating raw footage into a coherent story based on the plot scripts is a practical need, but it cannot be satisfied by the current video grounding techniques developed from existing datasets.
@@
  • We are the first to incorporate long-form videos with long abstract paragraphs for video grounding. Comparison results show the unique advantages of our dataset and the efficacy of our baseline in multi-paragraph video grounding.
@@ 2.2. Narrative Video Understanding
Understanding visual content presented in narrative videos is an important area with many works (Simoes et al., 2016; Haurilet et al., 2016; Nagrani and Zisserman, 2018; Huang et al., 2018; Gu et al., 2018; Del Fabro and Böszörmenyi, 2013; Huang et al., 2020; Rao et al., 2020; Tapaswi et al., 2016; Jasani et al., 2019; Soldan et al., 2022; Xiong et al., 2019; Bain et al., 2020; Sun et al., 2022) proposed accordingly. Most of these prior works neglect to model and understand the content of narrative videos based on their high-level storylines while focusing on specific downstream applications, such as movie genre classification (Simoes et al., 2016), character identification (Haurilet et al., 2016; Nagrani and Zisserman, 2018; Huang et al., 2018), action localization (Gu et al., 2018), scene segmentation (Del Fabro and Böszörmenyi, 2013; Huang et al., 2020), and shot classification (Rao et al., 2020). In addition to that, some works have started pursuing story-level understanding in many different ways. For instance, Tapaswi et al.  (Tapaswi et al., 2016) proposed MovieQA dataset to comprehend movie stories by question-answering. MSA (Xiong et al., 2019), CMD (Bain et al., 2020), and SyMoN (Sun et al., 2022) datasets utilized synopses as language queries and formulated movie understanding as text-to-video retrieval. This line of research is closely related to ours. However, we focus on a more challenging video grounding task with long contexts and complex queries, in which the model should understand the long-range cross-modal dependencies so as to reason about the video grounding results at story level.
@@ 3.1. Problem Formulation
Considering video paragraph grounding (Bao et al., 2021) is limited to a multi-query version of short single sentence grounding, we introduce a more challenging setting to incorporate long abstract paragraphs as queries called Multi-Paragraph Video Grounding (MPVG). Specifically, given an untrimmed video $\mathcal{V}$ and $N$ consecutive paragraph queries $\mathcal{Q}=\{Q_{1},Q_{2},...,Q_{N}\}$ as input, the output should be $N$ temporal intervals $\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{N}\}$ corresponding to each of the paragraph queries, where $\mathcal{T}_{i}=\left(t_{s}^{(i)},t_{e}^{(i)}\right)$ indicates the starting timestamp $t_{s}^{(i)}$ and ending timestamp $t_{e}^{(i)}$ for the $i$-th paragraph query in the target video. In our dataset, the video $\mathcal{V}$ is an episode from a TV drama, and $\mathcal{Q}$ is the corresponding human-written synopsis that contains $N$ paragraphs, with $Q_{i}$ indicating the $i$-th paragraph in the synopsis $\mathcal{Q}$. Illustration of our MPVG is in Figure 2.
@@ 3.2. Data Collection and Annotation
We collect all the TV drama episodes from a leading online platform Tencent Video with official acknowledgement and permission. The plot synopsis for each episode of the TV dramas is scraped from a specialized TV review website [^1] URL: https://www.tvmao.com. All texts are translated into English using Tencent Cloud Translation for research purpose only. that contains lots of synopses of the most popular TV drama episodes written by professionals. Synopses that are too long or too short are discarded to ensure an adequate number of paragraphs in each synopsis. A total of 520 licensed and high-viewership TV dramas with textual synopses are finally selected to constitute our dataset, and we randomly sample several episodes from each selected TV drama to further annotate. Specifically, annotators are asked to read and understand the synopsis in advance. They then thoroughly watch the corresponding TV drama episode to determine the starting and ending time of the video content depicted by each synopsis paragraph.
Our data annotation pipeline is organized into multiple rounds to ensure the annotation quality. Specifically, all collected videos are first divided into numerous disjoint subsets of videos. In each annotation round, synopsis paragraphs for videos in one subset will be annotated with timestamps and each annotator is told to provide a score to indicate level of confidence in the annotated results. Afterwards, we first discard samples with low confidence as an initial cleanup, and then some of the remaining samples are selected to be manually checked in terms of quality. If the annotation quality is thought of as satisfactory, the annotation process will move on to another unlabeled subset of video data. Otherwise, the current batch of data would be re-annotated. The above procedures are repeated by tens of annotators until we finish the annotation of all candidate samples. For post-annotation assessment, we randomly select a proportion of the annotated data to be re-annotated by other annotators. Concretely, we calculate the temporal IoU (Intersection over Union) between the two results from different annotators, which reaches a value of about 85%. This assessment result is much better than those of other datasets, such as the ActivityNet (Krishna et al., 2017), where different annotators only achieve an agreement degree around 70%. The higher degree of agreement across different annotators in our dataset verifies the effectiveness of our designed pipeline for data annotation and quality control.
@@ 3.3. Data Statistics
Detailed comparison with other datasets. In Table 2, we compare our dataset with other existing datasets in detail. As suggested, our videos are much longer in duration than those of Charades-STA (Gao et al., 2017), ActivityNet-Captions (Krishna et al., 2017), DiDeMo (Anne Hendricks et al., 2017), TACoS (Regneri et al., 2013) and Ego4d-NLQ (Grauman et al., 2022). Although the average video duration in our dataset is shorter than that of MAD, our total duration of videos is more than twice that of MAD, showing that our dataset is at a larger scale. Furthermore, the duration of target segments in our dataset is significantly longer while the normalized target span is still short, making our target moments challenging to be localized. Note that some datasets like MAD have shorter normalized target span than ours, but their short and general queries bring the harmful and undesirable semantic ambiguity issue as mentioned before. In addition, our dataset is the first to incorporate paragraph queries, and the average number of words in each query is significantly larger than those of other datasets, which greatly reduces the semantic ambiguity of the queries. Moreover, our synopsis queries involve both abstract expressions and concrete descriptions, enabling the model to learn semantic concepts at more diverse abstraction levels.
Data splits. As shown by the statistics in Table 3, we carefully divide the entire data into three non-overlapping splits for training, validation, and testing. Our training, validation, and testing sets consist of 3,187, 400, and 400 videos, respectively. It is worth noting that each video is an episode from a TV drama. Additionally, we manually guarantee half of the videos for validation/testing are sourced from TV dramas that are not selected for the training set.
@@ 4.4. Model Training
(6) $\mathcal{L}_{loc}$ $=\frac{1}{N}\sum_{i=1}^{N}\left[\frac{1}{\lambda_{1}}\mathcal{L}_{l1}(\mathcal{\hat{T}}_{i},\mathcal{T}_{i})+\mathcal{L}_{GIoU}(\mathcal{\hat{T}}_{i},\mathcal{T}_{i})\right],$
(7) $\mathcal{L}_{att}$ $=-\frac{1}{N}\sum_{i=1}^{N}\text{log}\left(\sum_{j=1}^{T}m_{ij}\cdot a_{ij}\right)$
where $\mathcal{L}_{l1}$ and $\mathcal{L}_{GIoU}$ are L1 and GIoU (Rezatofighi et al., 2019) losses, respectively. $\mathcal{\hat{T}}_{i}$ is the predicted time span for the $i$-th query and $\mathcal{T}_{i}$ is the ground-truth. $\mathcal{L}_{att}$ is an attention loss on the global query features. The term $a_{ij}$ indicates the attention weights between the $i$-th global query feature and the $j$-th video feature, while $m_{ij}$ is an indicator that takes 1 if the $j$-th video feature is inside the ground-truth interval of the $i$-th query, and 0 otherwise. This loss explicitly encourages the model to learn higher attention weights between text queries and visual elements that are correlated. In total, the training loss is defined as the weighted sum of the above two losses as $\mathcal{L}=\lambda_{1}\mathcal{L}_{loc}+\lambda_{2}\mathcal{L}_{att}$, where $\lambda_{1}$ and $\lambda_{2}$ are the hyper-parameters to balance these two different kinds of losses.
@@ 5.1. Experimental Setup
Implementation details. Our proposed method is implemented by PyTorch. The pre-extracted SlowFast, CLIP, and OCR features are aligned at sequence dimension and concatenated at channel dimension as the video feature input. We use a pre-trained RoBERTa (Liu et al., 2019) model to extract OCR features at each timestamp. The loss weights are set as $\lambda_{1}=1,\lambda_{2}=0.2$. For data augmentation, we choose to randomly shuffle the order of paragraphs in the same synopsis by a probability of $p$ during training and $p=\text{max}(0,1-\frac{T_{i}}{T_{max}})$, where $T_{i}$ is the index of the current training epoch and $T_{max}$ is set to 20. We adopt a local window length $M$ of 25 for our video encoder. The number of layers for video encoder and query decoder are set as 2 and 3, respectively. Our model is trained on 4 NVIDIA Tesla V100 GPUs by Adam (Kingma and Ba, 2014) optimizer using a learning rate of 0.0001 and batch size of 16 for a total of 50 epochs within one day.
@@ Acknowledgements.
This work was supported partially by the NSFC (U21A20471, U22A2095, 62076260, 61772570), Guangdong Natural Science Funds Project (2023B1515040025, 2022B1111010002), Guangdong NSF for Distinguished Young Scholar (2022B1515020009), Guangdong Provincial Key Laboratory of Information Security Technology (2023B1212060026), Guangzhou Science and Technology Plan Project (202201011134) and GACC-Research Projects (2022HK135).